OpenStudy (anonymous):

A large smartphone company produces 15,000 per hour. How long will it take this company to produce 300,000 smartphone? A. 15h B. 20h C. 30h D. 40h

OpenStudy (anonymous):

@shadow13 @aminamoazzam @Blade_Lover @Mehek14

Mehek (mehek14):

what do you think?

Mehek (mehek14):

divide \(\dfrac{15000}{300000}\)

Mehek (mehek14):

what do you get?

OpenStudy (anonymous):

20

Mehek (mehek14):

that's your answer

Mehek (mehek14):

\(\dfrac{15000}{1}=\dfrac{300000}{20}\)

OpenStudy (anonymous):

Thanks for your help @Mehek14